OpenStudy (anonymous):

Factor the expression x^4+8x^2-20??

OpenStudy (anonymous):

x^4+8x^2-20 ..?

OpenStudy (anonymous):

Were doing some cubic stuff in class lol

OpenStudy (anonymous):

its x^4+8x^2-20 @Algebraic!

OpenStudy (anonymous):

substitute u =x^2

OpenStudy (mathstudent55):

Start wit a substitution to make it simpler: Let u = x^2, then u^2 = x^4. Now re3place the x's with the u's: u^2 + 8u - 20 Can you factor that one?

OpenStudy (anonymous):

@mathstudent55 Im not sure.... My teacher didnt really go over this so i have no idea. Would I pull out a 4?

OpenStudy (anonymous):

u^2 + 8u - 20 is a quadratic... you've factored a lot of those by now I think....

OpenStudy (anonymous):

what two number multiply to give -20 and add to give 8?

OpenStudy (anonymous):

@Algebraic! 10 and -2

OpenStudy (anonymous):

(u+10)(u-2) = (x^2 +10)(x^2 -2)

OpenStudy (anonymous):

(x^2-2) is a difference of squares.. so you can factor that ...

OpenStudy (anonymous):

@Algebraic! Okay so what would i do with the other equation?

OpenStudy (anonymous):

(x^2+10) ? no real roots...